Angelena files as a head of household. In 2018, she reported $54,100 of taxable income, including a $15,400 qualified dividend. What is her gross tax liability, rounded to the nearest whole dollar amount? (Use the tax rate schedules.)
Answer #1

Out of 54100 , 15400 is qualified dividend so remaining 38700 will be taxed at normal rate.

So tax on ordinary income = 1360 + (38700-13700)*12% = 4360

In qualified dividend is taxed at 0% if income upto 51700.

Hence excess (54100-51700) will be taxed at 15%

HeNCE tax on dividend = 2400*15% = 360

Hence total tax = 4360+360 = 4720